a) A corporate bond with RM1000 maturity value carries a 7.5% coupon rate. It currently
makes interest payments semi-annually.
(i) This 12-year bond currently sells for RM961.88. What is the rate of return on this
bond?
(ii) If the bond sold for RM1,030.32, what is the rate of return on this bond?
1
Expert's answer
2016-02-20T00:00:58-0500
a) RM1000 maturity value, 7.5% coupon rate, interest payments semi-annually.
(i) If this 12-year bond currently sells for RM961.88, then the rate of return on this
bond is: Rate of return = (C + P1 - P0)/P0 = (1,000*0.075/2*24 + 961.88 - 1,000)/1,000 = 86.2%
(ii) If the bond sold for RM1,030.32, then the rate of return on this bond is: Rate of return = (C + P1 - P0)/P0 = (1,000*0.075/2*24 + 1,030.32 - 1,000)/1,000 = 93%